A cylindrical gasoline storage tank on Mr. Edwards' farm needs to be painted. The tank is 8 feet long and has a diameter of 4 feet. If one gallon of paint covers 350 square feet, how many cans of paint will Mr. Edwards need?
Answer by lwsshak3(11628)

A cylindrical gasoline storage tank on Mr. Edwards' farm needs to be painted. The tank is 8 feet long and has a diameter of 4 feet. If one gallon of paint covers 350 square feet, how many cans of paint will Mr. Edwards need?
...
r=2 ft
h=8 ft
surface area of tank=area of the two ends + area of cylindrical surface
area of two ends=2*πr^2=2*π2^2=8π
area of cylindrical part of tank=circumference*height=2πr*8=32π
Total area to be painted=8π+32π=40π sq ft
Cans of paint needed =40π/350=.35